OpenStudy (anonymous):

can you guys please show me the steps toward verifying this identity: cos(-x)/1+sin(-x) = sec x + tan x

OpenStudy (anonymous):

Is it \(\large \frac{\cos(-x)}{1+\sin(-x)}?\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First you have to know that \(\cos(-x)=\cos(x)\text{ and } \sin(-x)=-\sin(x)\). \[{\cos(-x) \over 1+\sin(-x)}={\cos(x) \over 1-\sin(x)}={(1+\sin(x))\cos(x) \over 1-\sin^2(x)}={(1+\sin(x))\cos(x) \over \cos^2(x)}\] \[={1+\sin(x) \over \cos(x)}={1 \over \cos(x)}+{\sin(x) \over \cos(x)}=sex(x)+\tan(x).\]

OpenStudy (anonymous):

\(\large \sec(x)+\tan(x)\)*

OpenStudy (anonymous):

ohh okay! thank you so much! c:

OpenStudy (anonymous):

"c:" << nice! :D

OpenStudy (anonymous):

ahah c: thanks!

OpenStudy (anonymous):

You're welcome!